# VP-MEL: Visual Prompts Guided Multimodal Entity Linking

### v1

@@ 1 Introduction
Figure 1: Comparison between Multimodal Entity Linking (MEL) and Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. VP-MEL aims to eliminate the reliance on mention words in the MEL task. By introducing visual prompts, it enhances the role of image information and balances the two modalities.
Multimodal entity linking (MEL) (Moon et al., 2018) task aims to utilize both image and text information to disambiguate the mention and link it to the correct entity in the knowledge base(KB).
Most MEL works (Gan et al., 2021; Wang et al., 2022a; Dongjie and Huang, 2022; Luo et al., 2023; Zhang et al., 2023a; Xing et al., 2023; Shi et al., 2024) mainly focus on improving the interaction of multimodal information and achieve promising performance. Existing methods represent mentions in the form of mention words and typically assume that each mention is associated with a high-quality image, which results in two limitations for MEL.
@@
Image modality impurity: Compared to text modality, image modality contains more noise. As shown in Figure 1, only the person on the right is relevant, while the one on the left is noise. Incorrect use of image information can significantly impact the results. Most existing coarse-grained methods (Yang et al., 2023; Luo et al., 2023; Wang et al., 2022a) encode the entire image directly, making it difficult to eliminate noise interference. Song et al. (2024) attempts to enhance visual features by utilizing fine-grained image attributes, but it heavily depends on high-quality images and may increase model size and training costs. To utilize images more effectively, it is necessary for MEL to find a way to overcome noise interference in images and focus on regions related to the mention.
@@ 3 Dataset
To address the VP-MEL task, we construct a VPWiki dataset as follows.
@@ 4.2 Visual Encoder
$\mathrm{F^{l}}=v_{[CLS]}^{0}\in V_{M_{j}}^{l}$
${V_{M_{j}}^{G}}^{{}^{\prime}}=\mathrm{LN}\left(\mathrm{Concat}(F^{3},F^{10},F^{11},F^{12})\right)$
$V_{M_{j}}^{G}=\mathrm{MLP}\left({V_{M_{j}}^{G}}^{{}^{\prime}}\right)$
Then, hidden states from the output layer of encoder $V_{M_{j}}^{l}$ are passed through a fully connected layer, which also transforms the dimensions to $d_{v}$, yielding the local features of the image $V_{M_{j}}^{L}\in\mathbb{R}^{(n+1)\times d_{v}}$.
$V_{M_{j}}^{L}=\mathrm{FC}\left(V_{M_{j}}^{l}\right)$
@@ 4.3 Detective-VLM
Most VLMs (Liu et al., 2024; Zhu et al., 2024; Ye et al., 2023; Li et al., 2023) choose to use the CLIP visual encoder, which means they have a better ability to focus on the markers in images compared to other visual methods (Cai et al., 2024; Shtedritski et al., 2023). Therefore, we instruction fine-tune a VLM, designed to extract effective information from images. The VLM follows template designed below to further mine potential information from image $m_{v_{j}}$ and sentence $m_{s_{j}}$ of mention $M_{j}$, assisting in subsequent feature extraction.
@@
$\min_{\theta}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i})$
@@
$\mathcal{L}(f_{\theta}(x_{i}),y_{i})=-\sum_{t=1}^{T}\log P_{\theta}(y_{i}^{(t)}|x_{i})$
@@ 4.4 Textual Encoder
For the mention $M_{j}$, after concatenating mention sentence $m_{s_{j}}$ with $m_{w_{j}}$, they form the input sequence, with different parts separated by [CLS] and [SEP] tokens.
$I_{M_{j}}=\left[CLS\right]m_{w_{j}}\left[SEP\right]m_{s_{j}}\left[SEP\right]$
@@
The input sequence for entity $E_{i}$ consists of the entity name $e_{n_{i}}$ and entity attributes $e_{a_{i}}$, which can be represented as,
$I_{E_{i}}=\left[CLS\right]e_{n_{i}}\left[SEP\right]e_{a_{i}}\left[SEP\right]$
@@ 4.5 Multimodal Feature Interaction
$S_{T}^{G2G}=T_{M_{j}}^{G}\cdot T_{E_{i}}^{G}$
To further uncover fine-grained clues within local features, TFI applies attention mechanism to capture context vector from the local features $T_{M_{j}}^{L}$ and $T_{E_{i}}^{L}$, producing the global-to-local similarity score $S_{T}^{G2L}$ between the global feature $T_{E_{i}}^{G}$ and the context vector.
$S_{T}^{G2L}=\mathrm{TFI_{G2L}}(T_{E_{i}}^{G},T_{M_{j}}^{L},T_{E_{i}}^{L})$
@@
CMFI performs a fine-grained fusion of features across modalities. It integrates the visual and textual features to generate a new context vector, $h_{e}$:
$h_{e}=\mathrm{CMFI}(T_{E_{i}}^{G},V_{E_{i}}^{L})$
@@ 4.6 Contrastive Learning
Based on the three similarity scores $S_{V}$, $S_{T}$, and $S_{C}$, the model is trained using contrastive loss function. For a mention $M$ and entity $E$, the combined similarity score is the average of the similarity scores from the three independent units.
$S(M,E)=(S_{V}+S_{T}+S_{C})/3$
@@
$\mathcal{L}_{O}=-log\frac{\exp(S(M_{j},E_{i}))}{\sum_{i}\exp(S(M_{j},E_{i}^{{}^{\prime}}))}$
@@
$\mathcal{L}_{X}=-log\frac{\exp(S_{X}(M_{j},E_{i}))}{\sum_{i}\exp(S_{X}(M_{j},E_{i}^{{}^{\prime}}))},X\in\{V,T,C\}$
@@
$\mathcal{L}=\mathcal{L}_{O}+\lambda(\mathcal{L}_{V}+\mathcal{L}_{T}+\mathcal{L}_{C})$
@@ 5.2 Main Results
Table 2(b) presents the experimental comparison results between FBMEL and the MEL baselines. During the training process, the Detective-VLM is removed from FBMEL, and its output is replaced with mention words from WikiDiverse. The metrics of our method are second only to MIMIC and MELOV, with a difference of only 0.27%, 0.28%, and 0.07% compared to MIMIC. FBMEL increases the emphasis on image features while reducing the influence of text features. It is reasonable that FBMEL performs slightly worse than MEL methods in the MEL task. Furthermore, the absence of visual prompts in images somewhat limits FBMEL’s ability to fully utilize visual information.
@@ 6 Conclusion
We introduce a new VP-MEL task and construct a dataset for the VP-MEL task and propose a FBMEL framework, there are still some limitations in this work.
VP-MEL expands the application scenarios of MEL, allowing users to directly annotate areas of interest within images. However, this requires the image and text to be correlated to some extent. When the image and text are not correlated, the performance of VP-MEL can be limited. We choose to use grounding box as the format for visual prompts. In real-world application scenarios, users may use any irregular shape to make markings. In future work, the design of visual prompts will be enhanced. We hope this work can encourage more research to apply the recent advanced techniques from both the NLP and computer vision communities to improve its performance.
@@ A.5 Feature Interaction Formula
$\overline{h}_{p}=\mathrm{MeanPooling}(V_{E_{i}}^{L})$
$h_{vc}=\mathrm{FC}(\mathrm{LayerNorm}(\overline{h}_{p}+V_{M_{j}}^{G}))$
$h_{vg}=\mathrm{Tanh}(\mathrm{FC}(h_{vc}))$
$h_{v}=\mathrm{LayerNorm}(h_{vg}*h_{vc}+V_{E_{i}}^{G})$
$S_{V}^{M2E}=h_{v}\cdot V_{M_{j}}^{G}$
@@
$Q,K,V=T_{E_{i}}^{L}W_{tq},T_{M_{j}}^{L}W_{tk},T_{M_{j}}^{L}W_{tv}$
$H_{t}=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{d_{T}}})V$
@@
$h_{t}=\mathrm{LayerNorm}(\mathrm{MeanPooling}(H_{t}))$
$S_{T}^{G2L}=\mathrm{FC}(T_{E_{i}}^{G})\cdot h_{t}$
@@
$h_{et},h_{mt}=\mathrm{FC}_{c1}(T_{E_{i}}^{G}),\mathrm{FC}_{c1}(T_{M_{j}}^{G})$
$H_{ev},H_{mv}=\mathrm{FC}_{c2}(V_{E_{i}}^{L}),\mathrm{FC}_{c2}(V_{M_{j}}^{L})$
@@
$\alpha_{i}=\frac{exp(h_{et}\cdot H_{ev}^{i})}{\sum_{1}^{n+1}exp(h_{et}\cdot H_{ev}^{i})}$
$h_{ec}=\sum\limits_{i}^{n}\alpha_{i}*H_{ev}^{i},i\in[1,2,...,(n+1)]$
$h_{eg}=\mathrm{Tanh}(\mathrm{FC}_{c3}(h_{et})$

### v2

@@ 1 Introduction
Figure 1: Comparison between MEL and VP-MEL task. In the absence of mention words, MEL faces challenges in performing entity linking for the image content. VP-MEL overcomes this limitation by utilizing visual prompts to link specific regions in the images to the correct entities in knowledge base.
Multimodal entity linking (MEL) (Moon et al., 2018) task aims to utilize both image and text information to disambiguate the mention and link it to the correct entity in the knowledge base(KB). Most MEL works (Gan et al., 2021; Wang et al., 2022a; Dongjie and Huang, 2022; Luo et al., 2023; Zhang et al., 2023a; Xing et al., 2023; Shi et al., 2024) mainly focus on improving the interaction of multimodal information and achieve promising performance. Existing methods represent mentions in the form of mention words and typically assume that each mention is associated with a high-quality image, which results in two limitations for MEL.
@@
Image modality impurity: Compared to text modality, image modality contains more noise. As shown in Figure 1, only the person on the right is relevant, while the one on the left is noise. Incorrect use of image information can significantly impact the results. Most existing coarse-grained methods (Yang et al., 2023; Luo et al., 2023; Wang et al., 2022a) encode the entire image directly, making it difficult to eliminate noise interference. Song et al. (2024) attempts to enhance visual features by utilizing fine-grained image attributes, but it heavily depends on high-quality images and may increase model size and training costs. To utilize images more effectively, it is necessary for MEL to find a way to overcome noise interference in images and focus on the regions related to the mention.
@@ 3 Dataset
To address the VP-MEL task, we construct a VPWiki dataset as follows:
@@ 4.2 Visual Encoder
$\mathrm{F^{l}}=v_{[CLS]}^{0}\in V_{M_{j}}^{l},$
${V_{M_{j}}^{G}}^{{}^{\prime}}=\mathrm{LN}\left(\mathrm{Concat}(F^{3},F^{10},F^{11},F^{12})\right),$
$V_{M_{j}}^{G}=\mathrm{MLP}\left({V_{M_{j}}^{G}}^{{}^{\prime}}\right).$
Then, hidden states from the output layer of encoder $V_{M_{j}}^{l}$ are passed through a fully connected layer, which also transforms the dimensions to $d_{v}$, yielding the local features of the image $V_{M_{j}}^{L}\in\mathbb{R}^{(n+1)\times d_{v}}$:
$V_{M_{j}}^{L}=\mathrm{FC}\left(V_{M_{j}}^{l}\right).$
@@ 4.3 Detective-VLM
Most VLMs (Liu et al., 2024; Zhu et al., 2024; Ye et al., 2023; Li et al., 2023) choose to use the CLIP visual encoder, which means they have a better ability to focus on the markers in images compared to other visual methods (Cai et al., 2024; Shtedritski et al., 2023). Therefore, we instruction fine-tune a VLM, designed to extract effective information from images. The VLM follows template designed below to further mine potential information from image $m_{v_{j}}$ and sentence $m_{s_{j}}$ of mention $M_{j}$, assisting in subsequent feature extraction:
@@
$\min_{\theta}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i}),$
@@
$\mathcal{L}(f_{\theta}(x_{i}),y_{i})=-\sum_{t=1}^{T}\log P_{\theta}(y_{i}^{(t)}|x_{i}),$
@@ 4.4 Textual Encoder
For the mention $M_{j}$, after concatenating mention sentence $m_{s_{j}}$ with $m_{w_{j}}$, they form the input sequence, with different parts separated by [CLS] and [SEP] tokens:
$I_{M_{j}}=\left[CLS\right]m_{w_{j}}\left[SEP\right]m_{s_{j}}\left[SEP\right].$
@@
The input sequence for entity $E_{i}$ consists of the entity name $e_{n_{i}}$ and entity attributes $e_{a_{i}}$, which can be represented as:
$I_{E_{i}}=\left[CLS\right]e_{n_{i}}\left[SEP\right]e_{a_{i}}\left[SEP\right].$
@@ 4.5 Multimodal Feature Interaction
$S_{T}^{G2G}=T_{M_{j}}^{G}\cdot T_{E_{i}}^{G}.$
To further uncover fine-grained clues within local features, TFI applies attention mechanism to capture context vector from the local features $T_{M_{j}}^{L}$ and $T_{E_{i}}^{L}$, producing the global-to-local similarity score $S_{T}^{G2L}$ between the global feature $T_{E_{i}}^{G}$ and the context vector:
$S_{T}^{G2L}=\mathrm{TFI_{G2L}}(T_{E_{i}}^{G},T_{M_{j}}^{L},T_{E_{i}}^{L}).$
@@
CMFI performs a fine-grained fusion of features across modalities. It integrates visual and textual features to generate a new context vector, $h_{e}$:
$h_{e}=\mathrm{CMFI}(T_{E_{i}}^{G},V_{E_{i}}^{L}).$
@@ 4.6 Contrastive Learning
Based on the three similarity scores $S_{V}$, $S_{T}$, and $S_{C}$, the model is trained using contrastive loss function. For a mention $M$ and entity $E$, the combined similarity score is the average of the similarity scores from the three independent units:
$S(M,E)=(S_{V}+S_{T}+S_{C})/3.$
@@
$\mathcal{L}_{O}=-log\frac{\exp(S(M_{j},E_{i}))}{\sum_{i}\exp(S(M_{j},E_{i}^{{}^{\prime}}))},$
@@
$\mathcal{L}_{X}=-log\frac{\exp(S_{X}(M_{j},E_{i}))}{\sum_{i}\exp(S_{X}(M_{j},E_{i}^{{}^{\prime}}))},X\in\{V,T,C\}.$
@@
$\mathcal{L}=\mathcal{L}_{O}+\lambda(\mathcal{L}_{V}+\mathcal{L}_{T}+\mathcal{L}_{C}),$
@@ 5.2 Main Results
Table 2(b) presents the experimental comparison results between FBMEL and the MEL baselines. During the training process, the Detective-VLM is removed from FBMEL, and its output is replaced with mention words from WikiDiverse. The metrics of our method are second only to MIMIC and MELOV, with a difference of only 0.27%, 0.28%, and 0.07% compared to MIMIC. FBMEL increases the emphasis on image features while reducing the influence of text features. It is reasonable that FBMEL performs slightly worse than MEL methods in the MEL task. Furthermore, the absence of visual prompts in images somewhat limits the ability of FBMEL to fully utilize visual information.
@@ 6 Conclusion
VP-MEL expands the application scenarios of MEL, allowing users to directly annotate areas of interest within images. However, this requires the image and text to be correlated to some extent. When the image and text are not correlated, the performance of VP-MEL can be limited. We choose to use grounding box as the format for visual prompts. In real-world application scenarios, users may use any irregular shape to make markings. In future work, the design of visual prompts will be enhanced. We hope this work can encourage more research to apply the recent advanced techniques from both natural language processing and computer vision fields to improve its performance.
@@ A.5 Feature Interaction Formula
$\overline{h}_{p}=\mathrm{MeanPooling}(V_{E_{i}}^{L}),$
$h_{vc}=\mathrm{FC}(\mathrm{LayerNorm}(\overline{h}_{p}+V_{M_{j}}^{G})),$
$h_{vg}=\mathrm{Tanh}(\mathrm{FC}(h_{vc})),$
$h_{v}=\mathrm{LayerNorm}(h_{vg}*h_{vc}+V_{E_{i}}^{G}),$
$S_{V}^{M2E}=h_{v}\cdot V_{M_{j}}^{G}.$
@@
$Q,K,V=T_{E_{i}}^{L}W_{tq},T_{M_{j}}^{L}W_{tk},T_{M_{j}}^{L}W_{tv},$
$H_{t}=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{d_{T}}})V,$
@@
$h_{t}=\mathrm{LayerNorm}(\mathrm{MeanPooling}(H_{t})),$
$S_{T}^{G2L}=\mathrm{FC}(T_{E_{i}}^{G})\cdot h_{t}.$
@@
$h_{et},h_{mt}=\mathrm{FC}_{c1}(T_{E_{i}}^{G}),\mathrm{FC}_{c1}(T_{M_{j}}^{G}),$
$H_{ev},H_{mv}=\mathrm{FC}_{c2}(V_{E_{i}}^{L}),\mathrm{FC}_{c2}(V_{M_{j}}^{L}),$
@@
$\alpha_{i}=\frac{exp(h_{et}\cdot H_{ev}^{i})}{\sum_{1}^{n+1}exp(h_{et}\cdot H_{ev}^{i})},$
$h_{ec}=\sum\limits_{i}^{n}\alpha_{i}*H_{ev}^{i},i\in[1,2,...,(n+1)],$
$h_{eg}=\mathrm{Tanh}(\mathrm{FC}_{c3}(h_{et}),$